News

Copernicus Sentinel-2A production issue on 11 February

11 February 2021

As a consequence of a wrong calibration parameter in the ground segment activated in the night between 10 and 11 February, the Copernicus Sentinel-2A L1C and L2A production systematically failed up to the roll-back to the previous ground segment configuration finalised in mid-morning today. The impact is from Sentinel-2A sensing orbit 29457 to 29461.

The recovery of the missing products will be started soon after the completion of the ongoing recovery (as announced on 9 February) and it is expected to be completed at the beginning of next week.

Navigation Menu